# Dynamics of the general $Q$-tensor model interacting with a rigid body

### v1

@@ 6. Maximal regularity of the linearized problem and exponential stability
$\mathrm{X}_{1}\coloneqq\{z\in\mathrm{Y}_{1}:v=0,\enspace\partial_{\nu}Q=0,\enspace\text{on}\enspace\partial\mathcal{O},\enspace\text{and}\enspace v=\ell+\omega\times y,\enspace\partial_{\nu}Q=0,\enspace\text{on}\enspace\partial\mathcal{S}_{0}\}$
@@
(6.3) $\tilde{\mathbb{E}}_{T}\coloneqq\mathbb{E}_{T}\times\mathbb{E}_{T}^{p},\enspace\text{where}\enspace\mathbb{E}_{T}^{p}\coloneqq\mathrm{L}^{p}\bigl{(}J;\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})\bigr{)},$
and $\|\cdot\|_{\mathbb{E}_{T}^{p}}$ represents the norm associated to $\mathbb{E}_{T}^{p}$. With regard to the stress tensor part in the surface integral, we include precisely its linear components to establish estimates in the proof of Theorem 4.2 in Section 8. As a result, in addition to the usual Cauchy stress tensor, we consider $-\nicefrac{{2\xi}}{{3}}(\Delta Q-Q)$, i. e.,
(6.4) $\mathrm{T}(v,p,Q)\coloneqq\nabla v+(\nabla v)^{\top}-p\operatorname{Id}_{3}-\frac{2\xi}{3}(\Delta Q-Q).$
As the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity already requires the associated operator to be invertible, we introduce a shift. It does not pose any problems for the local well-posedness, but it is only critical for the global well-posedness. Thus, for $\mu\geq 0$ as well as $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, $(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{T}$ and $(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, recalling $S_{\xi}$ and $\tilde{S}_{\xi}$ from (3.2), we consider the linearization
(6.5) $\partial_{t}{v}+(\mu-\Delta){v}+\nabla{p}-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\enspace\operatorname{div}{v}=0,&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(\widehat{Q})\nabla v+(\mu-(\Delta-\operatorname{Id}_{3})){Q}&=f_{2},&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\ell})^{\prime}+\mu\ell+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,p,Q)n\,\mathrm{d}\Gamma&=f_{3},&&\enspace\text{in}\enspace(0,T),\\ J_{0}({\omega})^{\prime}+\mu\omega+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,p,Q)n\,\mathrm{d}\Gamma&=f_{4},&&\enspace\text{in}\enspace(0,T),\\ {v}={\ell}+{\omega}\times y,\enspace\text{on}\enspace(0,T)\times\partial\mathcal{S}_{0},\enspace{v}=0,&\enspace\text{on}\enspace(0,T)\times\partial\mathcal{O},\enspace\partial_{\nu}Q=0,&&\enspace\text{on}\enspace(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\enspace Q(0)=Q_{0},\enspace\ell(0)&=\ell_{0}\enspace\text{and}\enspace\omega(0)=\omega_{0}.\end{aligned}\right.$
Below, we assert the main result of this section on the maximal $\mathrm{L}^{p}$-regularity of (6.5), up to a shift for general $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$ and without shift in the case $\widehat{Q}=0$.
@@
Let $p\in(1,\infty)$, $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, $T\in(0,\infty]$, $(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{T}$ and $(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$.
@@
    There is $\mu_{0}\geq 0$ such that for all $\mu>\mu_{0}$, (6.5) admits a unique solution $(v,Q,\ell,\omega,p)\in\tilde{\mathbb{E}}_{T}$, and there exists $C_{\mathrm{MR}}>0$ such that $\|(v,Q,\ell,\omega,p)\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}\bigl{(}\|(f_{1},f_{2},f_{3},f_{4})\|_{\mathbb{F}_{T}}+\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}\bigr{)}$. The constant $C_{\mathrm{MR}}>0$ can be chosen independent of $T$ in the case of homogeneous initial values.
@@
    If $\widehat{Q}=0$, then the assertion from (a) is also valid without shift, that means, there exists a unique solution $(v,Q,\ell,\omega,p)\in\tilde{\mathbb{E}}_{T}$ to (6.5) for $\mu=0$. In particular, there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0})$, if $\mathrm{e}^{\eta t}(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{\infty}$, then there is a unique solution $\mathrm{e}^{\eta t}(v,Q,\ell,\omega,p)\in\tilde{\mathbb{E}}_{\infty}$ to (6.5) with $\mu=0$. Besides, there exists a constant $C_{\mathrm{MR}}>0$ such that
    $\quad\|\mathrm{e}^{\eta(\cdot)}v\|_{\mathbb{E}_{\infty}^{v}}+\|\mathrm{e}^{\eta(\cdot)}p\|_{\mathbb{E}_{\infty}^{p}}+\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathbb{E}_{\infty}^{Q}}+\|\mathrm{e}^{\eta(\cdot)}\ell\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\omega\|_{\mathrm{W}^{1,p}(0,\infty)}$
@@
(6.6) $\partial_{t}{v}-\Delta{v}+\nabla{p}-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\enspace\operatorname{div}{v}=0,&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(\widehat{Q})\nabla v-(\Delta-\operatorname{Id}_{3}){Q}&=f_{2},&&\enspace\text{in}\enspace(0,T)\times\mathcal{F}_{0},\\ {v}=0,\enspace\partial_{\nu}Q&=0,&&\enspace\text{on}\enspace(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\enspace Q(0)&=Q_{0},&&\enspace\text{for}\enspace y\in\mathcal{F}_{0}.\end{aligned}\right.$
Following [21, Section 5], we introduce the operator corresponding to (6.6). First, we recall the Helmholtz projection $\mathbb{P}\colon\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\to\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ on $\mathrm{L}^{2}$ and invoke the Stokes operator $A_{2}$ on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ defined by
(6.7) $A_{2}v\coloneqq\mathbb{P}\Delta v,\enspace\text{for}\enspace v\in\mathrm{D}(A_{2})\coloneqq\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\cap\mathrm{H}_{0}^{1}(\mathcal{F}_{0})^{3}\cap\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0}).$
@@
$\Delta_{\mathrm{N}}^{1}Q\coloneqq(\Delta-\operatorname{Id}_{3})Q,\enspace\text{for}\enspace Q\in\mathrm{D}(\Delta_{\mathrm{N}}^{1})\coloneqq\{Q\in\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3}):\partial_{\nu}Q=0\enspace\text{on}\enspace\partial\mathcal{F}_{0}\}.$
@@
We collect auxiliary results for our analysis in the lemma below.
[…]
@@
    the operator $A_{\xi}(\widehat{Q})$ from (6.8) is sectorial on $\mathrm{X}_{0}^{\mathrm{lc}}$ with spectral angle $\phi_{A_{\xi}(\widehat{Q})}<\nicefrac{{\pi}}{{2}}$, and
@@ 6.1. Lifting operators and added mass
(6.9) $\tilde{\mathrm{T}}(v,p)\coloneqq\nabla v+(\nabla v)^{\top}-p\operatorname{Id}_{3}.$
A key concept is the added mass operator, see e. g. [16] or [18, Section 4.2] for further background. The most important aspects about this approach are the existence of lifting operators for the inhomogeneous interface condition and a representation formula of the pressure appearing in the surface integrals. This will be of central relevance for the reformulation of (6.5) in operator form. Let us consider
(6.10) $-\Delta w+\nabla\psi&=0,\enspace\operatorname{div}w=0,&&\enspace\text{for}\enspace y\in\mathcal{F}_{0},\\ w&=0,\enspace\text{for}\enspace y\in\partial\mathcal{O},\enspace w=\ell+\omega\times y,&&\enspace\text{for}\enspace y\in\partial\mathcal{S}_{0},\\ \int_{\mathcal{F}_{0}}\psi\,\mathrm{d}y&=0.\end{aligned}\right.$
The following result on the solvability of (6.10) is classical. We refer for instance to [26, Lemma 3.6].
[…]
@@
$-\Delta W_{i}+\nabla\Psi&=0,\enspace\operatorname{div}W_{i}=0,&&\enspace\text{for}\enspace y\in\mathcal{F}_{0},\\ W_{i}&=0,&&\enspace\text{for}\enspace y\in\partial\mathcal{O},\\ W_{i}=e_{i},\enspace\text{for}\enspace i=1,2,3\enspace\text{and}\enspace W_{i}&=e_{i-3}\times y,\enspace\text{for}\enspace i=4,5,6,&&\enspace\text{for}\enspace y\in\partial\mathcal{S}_{0},\end{aligned}\right.$
$\mathbb{B}_{i,j}=\int_{\mathcal{F}_{0}}DW_{i}:DW_{j}\,\mathrm{d}y$ is invertible, and for $\tilde{\mathrm{T}}$ from (6.9), we have $\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}\tilde{\mathrm{T}}(w,\psi)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times\tilde{\mathrm{T}}(w,\psi)n\,\mathrm{d}\Gamma\end{pmatrix}=\mathbb{B}\binom{\ell}{\omega}$.
@@
(6.11) $D_{\mathrm{fl}}(\ell,\omega)\coloneqq w\enspace\text{and}\enspace D_{\mathrm{pr}}(\ell,\omega)\coloneqq\psi.$
@@
(6.12) $\Delta\varphi=0,\enspace\text{in}\enspace\mathcal{F}_{0},\enspace\partial_{\nu}\varphi=h,\enspace\text{on}\enspace\partial\mathcal{F}_{0}.$
@@
The above lifting operators allow us to rewrite the resolvent problem linked to $v$ in terms of $\mathbb{P}v$ and $(\operatorname{Id}_{3}-\mathbb{P})v$ for the Helmholtz projection $\mathbb{P}$ on $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. This splitting idea for fluid-structure interaction problems was e. g. used by Raymond [34]. For a proof of the next lemma, we refer to [26, Proposition 3.7].
[…]
Assume $(f_{1},\ell,\omega)\in\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathbb{C}^{3}\times\mathbb{C}^{3}$, and recall the Stokes operator on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ from (6.7), the Dirichlet lifting operator $D_{\mathrm{fl}}$ from (6.11), the lifting operator $N$ associated to the Neumann problem (6.12) as well as $N_{\mathcal{S}_{0}}$ from (6.13). Then $(v,p)\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ satisfies
$\lambda v-\Delta v+\nabla p&=f_{1},\enspace\operatorname{div}v=0,&&\enspace\text{for}\enspace y\in\mathcal{F}_{0},\\ v=0,\enspace\text{for}\enspace y\in\partial\mathcal{O},\enspace v&=\ell+\omega\times y,&&\enspace\text{for}\enspace y\in\partial\mathcal{S}_{0},\end{aligned}\quad\text{if and only if}\right.$
[…]
@@
(6.15) $\lambda v-\Delta v+\nabla p-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\enspace\operatorname{div}v=0,&&\enspace\text{in}\enspace\mathcal{F}_{0},\\ \lambda Q+\tilde{S}_{\xi}(\widehat{Q})\nabla v-(\Delta-\operatorname{Id})Q&=f_{2},&&\enspace\text{in}\enspace\mathcal{F}_{0},\\ \lambda m_{\mathcal{S}}\ell+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,p,Q)n\,\mathrm{d}\Gamma&=f_{3},\enspace\lambda J_{0}\omega+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,p,Q)n\,\mathrm{d}\Gamma=f_{4},\\ v=0,\enspace\text{on}\enspace\partial\mathcal{O},\enspace v&=\ell+\omega\times y,\enspace\text{on}\enspace\partial\mathcal{S}_{0},\enspace\partial_{\nu}Q=0,&&\enspace\text{on}\enspace\partial\mathcal{F}_{0}.\end{aligned}\right.$
@@
(6.16) $K$ $=\begin{pmatrix}m_{\mathcal{S}}\operatorname{Id}_{3}&0\\ 0&J_{0}\end{pmatrix}+M,\enspace\text{with}\enspace M\binom{\ell}{\omega}=\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n)n\,\mathrm{d}\Gamma\end{pmatrix},$

### v2

@@ 6. Maximal regularity of the linearized problem and exponential stability
$\mathrm{X}_{1}\coloneqq\{z\in\mathrm{Y}_{1}:v=0,\kern 5.0pt\partial_{n}Q=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0pt\text{and}\kern 5.0ptv=\ell+\omega\times y,\kern 5.0pt\partial_{n}Q=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0}\}$
@@
(6.3) $\tilde{\mathbb{E}}_{T}\coloneqq\mathbb{E}_{T}\times\mathbb{E}_{T}^{\pi},\kern 5.0pt\text{where}\kern 5.0pt\mathbb{E}_{T}^{\pi}\coloneqq\mathrm{L}^{p}\bigl(J;\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})\bigr),$
and $\|\cdot\|_{\mathbb{E}_{T}^{\pi}}$ represents the norm associated to $\mathbb{E}_{T}^{\pi}$. With regard to the stress tensor part in the surface integral, we include precisely its linear components to establish estimates in the proof of Theorem 4.2 in Section 8. In addition to the usual Cauchy stress tensor, we thus incorporate $\nicefrac{{2\xi}}{{3}}(\Delta Q-Q)$, i. e.,
(6.4) $\mathrm{T}(v,\pi,Q)\coloneqq\nabla v+(\nabla v)^{\top}-\pi\operatorname{Id}_{3}-\frac{2\xi}{3}(\Delta Q-Q).$
We need to include the $Q$-terms in order to obtain the exponential stability for the linearization at zero, and to establish the estimates of the surface integrals involving the stress tensor.
As the maximal $\mathrm{L}^{p}(\mathbb{R}_{+})$-regularity already requires the associated operator to be invertible, we introduce a shift, i. e., instead of considering the original linearized problem, we add an artificial term of the form $\mu\operatorname{Id}$ on the left-hand side of the linearized problem, see (6.5) below. It does not pose any problems for the local well-posedness, but it is only critical for the global well-posedness. Thus, for $\mu\geq 0$ as well as $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, $(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{T}$ and $(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$, recalling $S_{\xi}$ and $\tilde{S}_{\xi}$ from (3.2), we consider the linearization
(6.5) $\left\{\begin{aligned} \partial_{t}{v}+(\mu-\Delta){v}+\nabla\pi-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\kern 5.0pt\operatorname{div}{v}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(\widehat{Q})\nabla v+(\mu-(\Delta-\operatorname{Id}_{3})){Q}&=f_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ m_{\mathcal{S}}({\ell})^{\prime}+\mu\ell+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{3},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ J_{0}({\omega})^{\prime}+\mu\omega+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{4},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T),\\ {v}={\ell}+{\omega}\times y,\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{S}_{0},\kern 5.0pt{v}=0,&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{O},\kern 5.0pt\partial_{n}Q=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\kern 5.0ptQ(0)=Q_{0},\kern 5.0pt\ell(0)&=\ell_{0}\kern 5.0pt\text{and}\kern 5.0pt\omega(0)=\omega_{0}.\end{aligned}\right$.
Below, we assert the main result of this section on the maximal $\mathrm{L}^{p}$-regularity of (6.5), up to a shift by $\mu>0$ for general $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$ and without shift in the case $\widehat{Q}=0$. Even though we only require this result for $p>4$, we provide the more general case $p\in(1,\infty)\setminus\{\nicefrac{{4}}{{3}}\}$ here. We do not need a more restrictive condition on $p$, because we consider a linearized problem and take into account $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$.
@@
Let $p\in(1,\infty)\setminus\{\nicefrac{{4}}{{3}}\}$, $\widehat{Q}\in\mathrm{C}^{1}(\overline{\mathcal{F}_{0}},\mathbb{S}_{0}^{3})$, $T\in(0,\infty]$, $(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{T}$ as well as $(v_{0},Q_{0},\ell_{0},\omega_{0})\in\mathrm{X}_{\gamma}$.
@@
    There is $\mu_{0}\geq 0$ such that for all $\mu>\mu_{0}$, (6.5) admits a unique solution $(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{T}$, and there exists $C_{\mathrm{MR}}>0$ such that $\|(v,Q,\ell,\omega,\pi)\|_{\tilde{\mathbb{E}}_{T}}\leq C_{\mathrm{MR}}\bigl(\|(f_{1},f_{2},f_{3},f_{4})\|_{\mathbb{F}_{T}}+\|(v_{0},Q_{0},\ell_{0},\omega_{0})\|_{\mathrm{X}_{\gamma}}\bigr)$. The constant $C_{\mathrm{MR}}>0$ can be chosen independent of $T$ in the case of homogeneous initial values.
@@
    If $\widehat{Q}=0$, then the assertion from (a) is also valid without shift, that means, there exists a unique solution $(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{T}$ to (6.5) for $\mu=0$. In particular, there exists $\eta_{0}>0$ such that for every $\eta\in[0,\eta_{0})$, if $\mathrm{e}^{\eta t}(f_{1},f_{2},f_{3},f_{4})\in\mathbb{F}_{\infty}$, then there is a unique solution $\mathrm{e}^{\eta t}(v,Q,\ell,\omega,\pi)\in\tilde{\mathbb{E}}_{\infty}$ to (6.5) with $\mu=0$. Besides, there exists a constant $C_{\mathrm{MR}}>0$ such that
    $\quad\|\mathrm{e}^{\eta(\cdot)}v\|_{\mathbb{E}_{\infty}^{v}}+\|\mathrm{e}^{\eta(\cdot)}\pi\|_{\mathbb{E}_{\infty}^{\pi}}+\|\mathrm{e}^{\eta(\cdot)}Q\|_{\mathbb{E}_{\infty}^{Q}}+\|\mathrm{e}^{\eta(\cdot)}\ell\|_{\mathrm{W}^{1,p}(0,\infty)}+\|\mathrm{e}^{\eta(\cdot)}\omega\|_{\mathrm{W}^{1,p}(0,\infty)}$
@@
(6.6) $\left\{\begin{aligned} \partial_{t}{v}-\Delta{v}+\nabla\pi-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\kern 5.0pt\operatorname{div}{v}=0,&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {\partial_{t}{Q}}+\tilde{S}_{\xi}(\widehat{Q})\nabla v-(\Delta-\operatorname{Id}_{3}){Q}&=f_{2},&&\kern 5.0pt\text{in}\kern 5.0pt(0,T)\times\mathcal{F}_{0},\\ {v}=0,\kern 5.0pt\partial_{n}Q&=0,&&\kern 5.0pt\text{on}\kern 5.0pt(0,T)\times\partial\mathcal{F}_{0},\\ v(0)=v_{0},\kern 5.0ptQ(0)&=Q_{0},&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0}.\end{aligned}\right$.
Following [22, Section 5], we introduce the operator corresponding to (6.6). First, we recall the Helmholtz projection $\mathbb{P}\colon\mathrm{L}^{2}(\mathcal{F}_{0})^{3}\to\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ on $\mathrm{L}^{2}$ and invoke the Stokes operator $A_{2}$ on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ defined by
(6.7) $A_{2}v\coloneqq\mathbb{P}\Delta v,\kern 5.0pt\text{for}\kern 5.0ptv\in\mathrm{D}(A_{2})\coloneqq\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\cap\mathrm{H}_{0}^{1}(\mathcal{F}_{0})^{3}\cap\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0}).$
@@
$\Delta_{\mathrm{N}}^{1}Q\coloneqq(\Delta-\operatorname{Id}_{3})Q,\kern 5.0pt\text{for}\kern 5.0ptQ\in\mathrm{D}(\Delta_{\mathrm{N}}^{1})\coloneqq\{Q\in\mathrm{H}^{3}(\mathcal{F}_{0},\mathbb{S}_{0}^{3}):\partial_{n}Q=0\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{F}_{0}\}.$
@@
We collect auxiliary results for our analysis in the lemma below. Part (a) will be an ingredient in the proof of the sectoriality of the operator associated with the linearized interaction problem, while part (b) will be used for the spectral analysis.
Lemma 6.3 ([22, Section 5]).
@@
[…]
@@ 6.1. Lifting operators and added mass
(6.9) $\tilde{\mathrm{T}}(v,\pi)\coloneqq\nabla v+(\nabla v)^{\top}-\pi\operatorname{Id}_{3}.$
A key concept is the added mass operator, see e. g. [17] or [19, Section 4.2] for further background. The most important aspects about this approach are the existence of lifting operators for the inhomogeneous interface condition and a representation formula of the pressure appearing in the surface integrals. This will be of central relevance for the reformulation of (6.5) in operator form. Let us consider
(6.10) $\left\{\begin{aligned} -\Delta w+\nabla\psi&=0,\kern 5.0pt\operatorname{div}w=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0},\\ w&=0,\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{O},\kern 5.0ptw=\ell+\omega\times y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{S}_{0},\\ \int_{\mathcal{F}_{0}}\psi\,\mathrm{d}y&=0.\end{aligned}\right$.
The following result on the solvability of (6.10) is classical. We refer for instance to [27, Lemma 3.6]. It provides an elementary representation of the solution to the Stokes problem with inhomogeneous boundary conditions at the interface, capturing the motion of the rigid body.
Lemma 6.4.
@@
$\left\{\begin{aligned} -\Delta W_{i}+\nabla\Psi&=0,\kern 5.0pt\operatorname{div}W_{i}=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0},\\ W_{i}&=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{O},\\ W_{i}=e_{i},\kern 5.0pt\text{for}\kern 5.0pti=1,2,3\kern 5.0pt\text{and}\kern 5.0ptW_{i}&=e_{i-3}\times y,\kern 5.0pt\text{for}\kern 5.0pti=4,5,6,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{S}_{0},\end{aligned}\right$.
$\mathbb{B}_{i,j}\coloneqq\int_{\mathcal{F}_{0}}DW_{i}:DW_{j}\,\mathrm{d}y$ is invertible, where $DW_{i}$ represents the gradient of $W_{i}$, and for $\tilde{\mathrm{T}}$ from (6.9), we have $\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}\tilde{\mathrm{T}}(w,\psi)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times\tilde{\mathrm{T}}(w,\psi)n\,\mathrm{d}\Gamma\end{pmatrix}=\mathbb{B}\binom{\ell}{\omega}$.
@@
(6.11) $D_{\mathrm{fl}}(\ell,\omega)\coloneqq w\kern 5.0pt\text{and}\kern 5.0ptD_{\mathrm{pr}}(\ell,\omega)\coloneqq\psi.$
@@
(6.12) $\Delta\varphi=0,\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\kern 5.0pt\partial_{n}\varphi=h,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{F}_{0}.$
@@
The above lifting operators allow us to rewrite the resolvent problem linked to $v$ in terms of $\mathbb{P}v$ and $(\operatorname{Id}_{3}-\mathbb{P})v$ for the Helmholtz projection $\mathbb{P}$ on $\mathrm{L}^{2}(\mathcal{F}_{0})^{3}$. Besides, we obtain a formula for the pressure in terms of $\mathbb{P}v$ as well as the rigid body velocities. This will be important for handling the pressure term appearing in the surface integrals of the stress tensor, and it paves the way for a reformulation in operator form. This splitting idea for fluid-structure interaction problems was e. g. used by Raymond [36]. For a proof of the next lemma, we refer to [27, Proposition 3.7].
Lemma 6.5.
Assume $(f_{1},\ell,\omega)\in\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})\times\mathbb{C}^{3}\times\mathbb{C}^{3}$, and recall the Stokes operator on $\mathrm{L}_{\sigma}^{2}(\mathcal{F}_{0})$ from (6.7), the Dirichlet lifting operator $D_{\mathrm{fl}}$ from (6.11), the lifting operator $N$ associated to the Neumann problem (6.12) as well as $N_{\mathcal{S}_{0}}$ from (6.13). Then $(v,\pi)\in\mathrm{H}^{2}(\mathcal{F}_{0})^{3}\times\mathrm{H}^{1}(\mathcal{F}_{0})\cap\mathrm{L}_{0}^{2}(\mathcal{F}_{0})$ satisfies
$\left\{\begin{aligned} \lambda v-\Delta v+\nabla\pi&=f_{1},\kern 5.0pt\operatorname{div}v=0,&&\kern 5.0pt\text{for}\kern 5.0pty\in\mathcal{F}_{0},\\ v=0,\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{O},\kern 5.0ptv&=\ell+\omega\times y,&&\kern 5.0pt\text{for}\kern 5.0pty\in\partial\mathcal{S}_{0},\end{aligned}\quad\text{if and only if}\right$.
(6.14) $\left\{\begin{aligned} \lambda\mathbb{P}v-A_{2}\mathbb{P}v+A_{2}\mathbb{P}D_{\mathrm{fl}}(\ell,\omega)&=\mathbb{P}f_{1},\\ (\operatorname{Id}_{3}-\mathbb{P})v&=(\operatorname{Id}_{3}-\mathbb{P})D_{\mathrm{fl}}(\ell,\omega),\\ \pi&=N(\Delta\mathbb{P}v\cdot n)-\lambda N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n).\end{aligned}\right$.
@@
(6.15) $\left\{\begin{aligned} \lambda v-\Delta v+\nabla\pi-\operatorname{div}S_{\xi}(\widehat{Q})(\Delta-\operatorname{Id}_{3})Q&=f_{1},\kern 5.0pt\operatorname{div}v=0,&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\\ \lambda Q+\tilde{S}_{\xi}(\widehat{Q})\nabla v-(\Delta-\operatorname{Id})Q&=f_{2},&&\kern 5.0pt\text{in}\kern 5.0pt\mathcal{F}_{0},\\ \lambda m_{\mathcal{S}}\ell+\int_{\partial\mathcal{S}_{0}}\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{3},\\ \lambda J_{0}\omega+\int_{\partial\mathcal{S}_{0}}y\times\mathrm{T}(v,\pi,Q)n\,\mathrm{d}\Gamma&=f_{4},\\ v=0,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{O},\kern 5.0ptv&=\ell+\omega\times y,\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{S}_{0},\kern 5.0pt\partial_{n}Q=0,&&\kern 5.0pt\text{on}\kern 5.0pt\partial\mathcal{F}_{0}.\end{aligned}\right$.
@@
(6.16) $K$ $=\begin{pmatrix}m_{\mathcal{S}}\operatorname{Id}_{3}&0\\ 0&J_{0}\end{pmatrix}+M,\kern 5.0pt\text{with}\kern 5.0ptM\binom{\ell}{\omega}=\begin{pmatrix}\int_{\partial\mathcal{S}_{0}}N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n)n\,\mathrm{d}\Gamma\\ \int_{\partial\mathcal{S}_{0}}y\times N_{\mathcal{S}_{0}}((\ell+\omega\times y)\cdot n)n\,\mathrm{d}\Gamma\end{pmatrix}$,